OpenStudy (anonymous):

Line is parallel to y=-2/3x+3 and passes through (3,-4). Graph both lines.

OpenStudy (johnweldon1993):

remember a parallel line has the same slope as the other line

OpenStudy (anonymous):

let the other line be \(y=mx+b\) 1. parallel lines have identical slopes 2. passes through the given point. so, you can now get your y-intercept

OpenStudy (anonymous):

now that you have the complete equation, you can happily draw it

OpenStudy (johnweldon1993):

so now that you have the slope....and a point you want to pass through use the point slope formula y - y1 = m(x - x1) where y1 is your y point x1 is your x point and m is your slope

OpenStudy (anonymous):

by the way, just noticed that your oriinal line does not have the "x" term in it

OpenStudy (anonymous):

thanks guys.

OpenStudy (anonymous):

a medal is in order